OpenStudy (anonymous):

which rule would apply to the expression w^2w^3

OpenStudy (shane_b):

The exponent rule which states that:\[a^xa^y=a^{x+y}\]This is based on: \[(a*a) * (a*a*a)=a*a*a*a*a*a=a^6\]

OpenStudy (shane_b):

err...I missed an "a" in the second group :P

OpenStudy (unklerhaukus):

2+3=5≠6

OpenStudy (anonymous):

ok thanks for the help